# Bridge indices of spatial graphs and diagram colorings

### v1

@@ 5.4 Eulerian spatial graphs from links
First, consider the two-component spatial graph shown in Figure 17 (left). Note that this graph is almost unknotted and, perhaps surprisingly, has bridge index $3$ (which we will verify in Example 5.17). Further note that changing the number of twists in various places, as long as each twist region remains an odd number of half-twists, will not change the linkedness, almost unknottedness, or bridge index of the graph.
Second, consider the $\Theta_{4}$-graph shown in Figure 17 (right). Note that this graph is not almost unknotted, but it also has bridge index $3$. Further note that changing the number of twists in various places, as long as the parity of the number of half-twists of each twist region is preserved, will not change the bridge index of the graph.
For both graphs, if we create a family of graphs from them by adding further pairs of edges that are clasped together in the same way as the two pairs in the given graphs, the upper bound for the bridge index coming from our algorithm increases accordingly, and we conjecture that this upper bound is in fact the bridge index. Note however that the almost unknottedness of the spatial graph in Figure 17 (left) will be destroyed once more edges are added.
Fact 5.15.
Let $G_{n}$ be the $\Theta_{2n}$-graph obtained by increasing the number of edges of either of the graphs in Figure 17 as described above. Then the bridge index of $G_{n}$ is less than or equal to $n+1$.
@@
Another way of viewing the construction in Example 5.14 is by fusing pairs of maxima and minima of pretzel links, which leads us to a more general construction. Recall that a graph is Eulerian if every vertex has even degree. There are numerous collections of knots and links that have an arbitrarily large number of quandle colorings or admit Coxeter quotients with arbitrarily large reflection rank. Therefore, the knots in these collections have arbitrarily large bridge index. Taking a knot and fusing pairs of local maxima or pairs of local minima together creates an Eulerian graph with degree four vertices (see Figure 18). In various cases, we can obtain an almost unknotted graph this way. Furthermore, we inherit some lower bounds from the links.
Figure 18: Fusing two maxima to form a vertex of degree four.
Recall from Section 4.2 that we can use labelings by a Coxeter group to obtain lower bounds for bridge indices. Following [blair2024coxeter], we call a surjective group homomorphism $\rho\colon\pi_{1}(S^{3}\backslash L)\twoheadrightarrow C$ from the fundamental group of a link to a Coxeter group $C$ a good quotient if meridians are mapped to reflections.
@@
Suppose that $\rho$ is a good quotient for a link $L$. Let $D$ be a bridge index minimizing diagram of $L$. Let $G_{L}$ be a spatial graph represented by the result of fusing the two local extrema of $D$ with two different labels together (see Figure 19 for a demonstration). Then $\rho$ induces a surjective homomorphism $\pi_{1}(S^{3}\backslash G_{L})\twoheadrightarrow C$ sending meridians around the fusing vertex to a reflection.
@@
Take two local extrema in $D$ associated to reflections $x,y$. After fusing, we can label the strands connected to the newly created degree four vertex as $y$. The presentation of the Coxeter group that $\rho$ surjects to is changed from $\langle x_{1},x_{2},\cdots,x_{n-2},x,y\;\ |\;\ R,\;\ (xx_{i})^{w_{i}},(yx_{j})^{w_{j}}\rangle$ to $\langle x_{1},x_{2},\cdots,x_{n-2},y\;\ |\;\ R,\;\ (yx_{i})^{w_{i}},(yx_{j})^{w_{j}}\rangle$. Fusing the two extrema of $D$ changes the graph $\Delta$ by a collapsing move as shown in Figure 20. In particular, we set two generators equal to each other. Thus $\rho$ induces a surjective homomorphism from the fused spatial graph to a Coxeter group with one fewer reflection rank. ∎
@@
Figure 19: Fusing two maxima of a generalized pretzel link together. Before fusing, we have a good quotient to a Coxeter group of reflection rank 4. After fusing, we have a good quotient to a Coxeter group of reflection rank 3.
[…]
Figure 20: Fusing two maxima affects $\Delta$ (from Definition 4.5) by identifying two vertices and preserving adjacency. For instance, the reflection $x$ was connected to $w$ and $z$ beforehand. After fusing, we have that $y$ is connected to $w$ and $z$. The circles in the diagrams on the left hand side represent different rational tangles.
Example 5.17 (Graphs from Montesinos links).
Consider the left picture of Figure 21. Once rational tangles are added in the circles, this is a Montesinos link, and the authors of [baader2021coxeter] have shown that it admits a good quotient. In particular, the meridional rank and the bridge index is $n$, where $n$ is the number of rational tangles inserted. Each maximum is associated to a reflection. Using Proposition 5.16, we can obtain a spatial graph with bridge index $n-1$. Thus we can construct infinitely many graphs for which our technique can produce the exact bridge index, where in this case, the graphs are not just coming from a family of vertex sums as in previous examples. Of course, some choices will not give an almost unknotted graph, but some choices, such as Figure 17 (left), will.
Corollary 5.18.
Let $G_{n}$ be the spatial graph obtained from fusing the maxima and minima of a Montesinos link with $n$ rational tangles. Then the bridge index of $G_{n}$ is $n-1$.
Figure 21: A schematic to construct infinitely many graphs for which our technique can produce the exact bridge index. The circles represent different rational tangles.
[…]
Sarah Blackwell
University of Virginia
Email address: blackwell@virginia.edu
URL: https://seblackwell.com/
Puttipong Pongtanapaisan
Arizona State University
Email address: ppongtan@asu.edu
URL: https://puttipongtanapaisan.wordpress.com/
Hanh Vo
Arizona State University
Email address: thihanhv@asu.edu
URL: https://hanhv.github.io/

### v2

@@ 5.4 Eulerian spatial graphs from links
These examples are related to constructions in [FM17], where the authors were interested in nontriviality of the spatial graphs, but here we pay attention to the bridge index.
Example 5.16 (Bouquet graphs).
First we consider the family of almost unknotted bouquet graphs $G_{n}$ shown in Figure 25. Using the terminology in [FM17], the graph $G_{n}$ is a vertex sum of a Montesinos tangle, where no rational subtangle has the $\infty$ parity. This means that the arc in the diagram connected to the northwest endpoint of each tangle has the other endpoint in the southeast corner. By Theorem 3.3 of [FM17], the spatial graph is ravel, which means that the spatial graph itself is nontrivial but every cycle is unknotted. Since a bouquet graph has the two cycles as proper subgraphs, being ravel is the same as being almost unknotted.
We claim that the bridge index of $G_{n}$ is at most $n+1$. The standard diagram of a Montesinos knot containing $n$ rational tangles has $n$ local maxima. To form a bouquet graph, we perform a vertex sum, which fuses a local maximum with a local minimum (see Figure 28). This gives an embedding with n1n-1 local maxima and a vertex of degree four on one side of the bridge sphere. On the other side of the bridge sphere, there are $n+1$ local minima. Thus, the number of intersections of $G_{n}$ with the bridge sphere is $2(n+1).$ Dividing by $2$ gives the claim.
To show that the bridge index is at least $n+1$, we will use Proposition 4.13. The length $n$ Montesinos links admits a rank $n$ Coxeter quotient [BKM24]. Performing a vertex closure as indicated in the figure still yields a rank $n$ Coxeter quotient since the strands where the closure takes place are sent to the same reflection. Therefore, each bridge splitting must have $n$ components in one trivial tangle. Since one component has weight 4 and the other n1n-1 components have weight $2$, the bridge index is $\frac{2(n-1)+4}{2}=n+1.$
Corollary 5.17.
The bouquet graph $G_{n}$ in Figure 25 is almost unknotted and has bridge index $n+1.$
Figure 25: A family of almost unknotted bouquet graphs with $n$ rational subtangles drawn as dashed circles.
@@
First, consider the two-component spatial graph shown in Figure 26 (left). Note that this graph is almost unknotted and has unweighted bridge index $3$ (which we will verify in Example 5.20). Further note that increasing the number of twists in various places, as long as each twist region remains an odd number of half-twists, will not change the linkedness, almost unknottedness, or bridge index of the graph.
Second, consider the $\Theta_{4}$-graph shown in Figure 26 (right). Note that this graph is not almost unknotted, but it also has unweighted bridge index $3$. Further note that increasing the number of twists in various places, as long as the parity of the number of half-twists of each twist region is preserved, will not change the bridge index of the graph.
For both graphs, if we create a family of graphs from them by adding further pairs of edges that are clasped together in the same way as the two pairs in the given graphs, as shown in Figure 27 for Figure 26 (left), the upper bound for the bridge index coming from our algorithm increases accordingly, and we conjecture that this upper bound is in fact the bridge index. Note however that the almost unknottedness of the spatial graph in Figure 26 (left) will be destroyed once more edges are added.
Fact 5.19.
Let $G_{n}$ be the $\Theta_{2n}$-graph obtained by increasing the number of edges of either of the graphs in Figure 26 as described above. Then the unweighted bridge index $\widehat{\beta}(G_{n})$ is less than or equal to $n+1$, and the bridge index $\beta(G_{n})$ is less than or equal to $2n$.
@@
Figure 27: A family of graphs created from Figure 26 (left) by adding further pairs of clasped edges.
Another way of viewing the construction in Example 5.18 is by fusing pairs of maxima and minima of pretzel links, which leads us to a more general construction. Recall that a graph is Eulerian if every vertex has even degree. There are numerous collections of knots and links that have an arbitrarily large number of quandle colorings [Prz98] or admit Coxeter quotients with arbitrarily large reflection rank [BKM24]. Therefore, the knots in these collections have arbitrarily large bridge index. Taking a knot and fusing pairs of local maxima or pairs of local minima together creates an Eulerian graph with degree four vertices (see Figure 28). In various cases, we can obtain an almost unknotted graph this way. Furthermore, we inherit some lower bounds from the links.
Figure 28: Fusing two extrema to form a vertex of degree four.
Recall from Section 4.2 that we can use labelings by a Coxeter group to obtain lower bounds for bridge indices. Following [BKM24], we call a surjective group homomorphism $\rho\colon\pi_{1}(S^{3}\backslash L)\twoheadrightarrow C$ from the fundamental group of a link to a Coxeter group $C$ a good quotient if meridians are mapped to reflections.
Example 5.20 (Graphs from Montesinos links).
Consider the left picture of Figure 29. Once rational tangles are added in the circles, this is a Montesinos link, and the authors of [BBK21] have shown that it admits a good quotient. In particular, the meridional rank and the bridge index is $n$, where $n$ is the number of rational tangles inserted. Each maximum is associated to a reflection. The right of Figure 29 gives a bridge position of the spatial graph of interest with n1n-1 maxima, so the unweighted bridge index is at most n1n-1. Suppose that we have another bridge position for the graph with $b$ maxima. We can replace the vertices with the trivial $2$-tangles that recover the original Montesinos link. This is then a bridge position for the link with $b+1$ maxima. By [BBK21], we have $b+1\geq n,$ so $b\geq n-1$, the unweighted bridge index is equal to n1n-1, and the bridge index is equal to $n$.
Thus we can construct infinitely many graphs for which our technique can produce the exact bridge index, where in this case, the graphs are not just coming from a family of vertex sums as in previous examples. Of course, some choices will not give an almost unknotted graph, but some choices, such as Figure 26 (left), will.
Corollary 5.21.
Let $G_{n}$ be the spatial graph obtained from fusing some of the maxima and minima of a Montesinos link with $n$ rational tangles, as depicted in Figure 29. Then the unweighted bridge index $\widehat{\beta}(G_{n})$ equals n1n-1, and the bridge index $\beta(G_{n})$ equals $n.$
Figure 29: A schematic to construct infinitely many graphs for which our technique can produce the exact bridge index. The dashed circles represent different rational tangles.
We now prove a general result on lower bounds of bridge index from repeatedly replacing vertices with rational tangles. In particular, we can use this to recover Corollary 5.21, but it is also useful in other situations.
@@
Suppose that $v$ is a vertex of a spatial graph $G$ of even degree $d$. If we replace $v$ with a $\tfrac{d}{2}$-string trivial tangle to get a spatial graph $G^{\prime}$ with one fewer vertex, then there is a surjection from the fundamental group of the exterior of $G$ to that of $G^{\prime}$ taking meridians to meridians.
@@
The boundary of the exterior of $G$ is a surface $F$. The tangle replacement that replaces a small neighborhood of a vertex with a $\tfrac{d}{2}$-trivial tangle can be thought of as attaching $\tfrac{d}{2}-1$ three-dimensional 2-handles with attaching regions on $F$; see Figure 30. By Van Kampen’s theorem, 2-handle additions only add relations to the fundamental group. In other words, the fundamental group of the exterior of $G^{\prime}$ is a quotient of the group for $G$. ∎
Figure 30: Replacing a vertex with rational tangles has the effect of attaching some number of 22-handles. In this particular figure, the vertex has degree six, and we attach 22-handles along the red and blue curves to get a $3$-string tangle.
5.5 Arbitrarily large bridge index
The goal of this subsection is to prove Theorem 1.2. We begin with a series of lemmas.
Lemma 5.23.
A diagram of a bouquet graph of $n$ petals with Wirtinger number $n$ coming from one seed is planar.
Proof.
The unique vertex $v$ of the graph has $2n$ edges emerging from it. If one seed extends to the entire diagram, then there is an embedding where one trivial tangle is a connected tree containing $v$. The other trivial tangle is made up of $n$ arcs containing local minima. A diagram representing this graph may contain crossings as we move away from $v$ traversing along edges. However, each crossing can be removed by the Reidemeister move (see move R6 from [KPW24], for example) that allows twisting near vertices until no crossings are left. ∎
As mentioned in [Jan+16, Tay21], it is not understood in general how clasping moves affect spatial graphs besides preserving almost unknottedness. In this paper, we show that by picking the clasping carefully in such a way that the quandle colorings are compatible, we can drive up the bridge index to be arbitrarily large while still being almost unknotted. This method can be modified to be compatible for colorings for even more general algebraic structures such as biquandles, but we just focus on the dihedral quandle for now.
Recall that a knot or spatial graph which admits a nontrivial quandle coloring by the dihedral quandle of order three $R_{3}$ is called tricolorable. For convenience, by a Type I replacement, we mean the replacement $T\mapsto T^{\prime}$ on three strands with the same color as shown on the right in Figure 31. By a Type II replacement, we mean the replacement $T\mapsto T^{\prime}$ where the colorings at $\partial T^{\prime}$ are as in Figure 32 (the colors are interchangeable, but the point is that the upper right and lower right endpoints of the tangle get one color and the remaining endpoints get another color).
Remark 5.24.
The conditions on the dihedral colorings at $\partial T^{\prime}$ are important. The reader can verify that for some other choice of colorings at $\partial T^{\prime}$, the number of colorings is strictly less than claimed in the next lemma.
Figure 31: A clasping move that increases the bridge index by one.
Figure 32: Three tricolorings of $T^{\prime}$ where there are two colors at the tangle boundaries.
Lemma 5.25.
Suppose $D$ is a diagram of a tricolorable almost unknotted graph with the number of colorings $Col_{R_{3}}(D)$. After one instance of the tangle replacement of Type I or Type II, we have that $D^{\prime}$ is still almost unknotted and $Col_{R_{3}}(D^{\prime})=3\cdot Col_{R_{3}}(D)$.
Proof.
To see that $D^{\prime}$ represents an almost unknotted graph, we view the tangle replacement in two stages as shown in Figure 31. The first stage comprises entirely of Reidemeister moves which do not change the spatial graph type. The second stage is the clasping move, which is known to preserve almost unknottedness.
The colorings shown in Figure 31 and Figure 32 also demonstrate the claim on the formula for the quandle coloring number of $D^{\prime}$. To be more precise, we name all the tricolored diagrams of $D$ as $D_{1},D_{2},\ldots,D_{Col_{R_{3}}(D)}$. After the replacement, we see that each colored diagram $D_{j}$ corresponds to 3 colorings of $D^{\prime}$.
Namely, for the Type I replacement, the three colorings are the version where all strands of $T^{\prime}$ are colored the same and two additional non-monochromatic colors for $T^{\prime}$ (in Figure 31, notice that green and blue can be swapped). For the Type II replacement, the three colorings for each $D_{j}$ are shown in Figure 32. In conclusion, the total number of colorings of $D^{\prime}$ is $3\cdot Col_{R_{3}}(D)$ as claimed. ∎
The next lemma is straightforward. We use the same notation for the tangle replacement $T\mapsto T^{\prime}$. In Figure 33, the opaque gray oval is shown to mark the spot where tangle replacements can be performed to drive up the bridge index.
Lemma 5.26.
[…]
Proof.
Maximally extend the $k$ seeds for $D\backslash T^{\prime}$. By assumption, the strands that intersect $\partial T=\partial T^{\prime}$ labeled with the letter $x$ are colored by the end of this process. Parts of these strands are contained in $T^{\prime},$ but the current choice of seeds is not enough to color the entire $T^{\prime}.$ An additional seed of weight $2$ (colored blue in Figure 33) can be used to complete the coloring to the entire tangle $T^{\prime}$. ∎
@@ 5.5 Arbitrarily large bridge index
Figure 33: Suppose that the diagram before the tangle replacement has Wirtinger number $k$ satisfying Lemma 5.26. Then, the diagram after the tangle replacement has Wirtinger number $k+1.$ Note that a Reidemeister 2 move has been performed to make the two seeds extend. We have put in an opaque gray oval to indicate the spot to iteratively insert $T^{\prime}$ to drive the bridge index to be as large as needed. The sequence of diagrams should be viewed left to right, top to bottom.
We are now ready to prove Theorem 1.2. We will repeatedly discuss $3$-string tangle replacements, and so it will be convenient to name endpoints of the tangle NE, NC, NW, SE, SC, and SW as depicted in Figure 34. Throughout, we are using Theorem 1.1 for upper bounds and Proposition 4.5 for lower bounds. Note that while Proposition 4.5 is a lower bound in terms of $\widehat{\beta}$ (the unweighted version), it still works for bouquet graphs as there is no ambiguity of where to place the vertex: the vertex has to be in one of the trivial tangles. In contrast, when one has a $\theta_{n}$-graph, for example, the choice of whether to put the vertices on the same side will affect the weighted versus the unweighted version.
@@
Figure 34: Extending the seed colorings through each twist knot.
See 1.2
Proof.
Observe that the bouquet graph of $n$ petals has Euler characteristic (n1)-(n-1). Thus, by taking $n$ to infinity, we obtain all possible Euler characteristics for interesting graph types that admit almost unknotted embeddings.
Let $K_{1},K_{2},\ldots,K_{n}$ be tricolorable twist knots. In particular, each $K_{i}$ is a closure of the tangle in Figure 34 and each of these admit an embedding with two local maxima and two local minima. We fuse the lower left local minima to become the bouquet graph vertex shown on the left of Figure 35. We also splice the lower right local minima in the manner shown in the figure, and call the resulting graph $B_{n}$.
Figure 35: An almost unknotted bouquet graph of $n$ petals which we will perform clasping to. The blue strand corresponds to a cycle. There are $n$ rectangular boxes each corresponding to a rational tangle defining $K_{n}$.
@@
The graph $B_{n}$ has bridge index precisely $n+1.$ To see this, note that the green and blue strands in a diagram $D$ as shown in Figure 36 function as the two seeds: one with weight $2n$ and one with weight $2$. The lower bound for this claim on $B_{n}$ comes from our assumption on tricolorable twist knots. As before, handling the lower bound, treat the colors at the seed to be quandle labels. This means that the quandle labels propagate consistently so that the NW and SW strands receive the same color, say green. Furthermore, the remaining SC and SW strands receive the same color, say blue.
@@
Now, to see that the bridge index can be arbitrarily large for a fixed $n,$ we start with $B_{n}$ and perform a tangle replacement $T\rightarrow T^{\prime}$ as shown in Figure 36, and call the resulting graph $B_{n}^{2}$, represented by a diagram $D^{2}$. By Lemma 5.25 and Lemma 5.26, $B^{2}_{n}$ has bridge index precisely $n+2$. For the upper bound, note that the original two seeds from $B_{n}$ suffice to extend the coloring of $D\backslash T^{\prime}$ maximally to satisfy the hypothesis of Lemma 5.26 (the NW, SW, SC, and SE strands are colored). Therefore, an additional seed of weight $2$ will suffice to generate the coloring for the entire diagram. For the lower bound, observe that this is a Type I replacement: the diagram $D$ is tricolorable, and we are performing a replacement near a vertex where all arcs are colored the same.
Finally, the process to construct further iterations $B_{n}^{3},B_{n}^{4},\ldots,B_{n}^{k}$ can be described as follows. Now that $B_{n}^{2}$ already contains an instance of $T^{\prime}$, we look inside $T^{\prime}$ to perform more tangle replacements sequentially at the opaque gray oval in Figure 33. We remark again that each replacement satisfies Lemma 5.25 and Lemma 5.26. For the upper bound, notice that from the first picture to the last of Figure 33, we can extend the colorings at the strands marked $x$ to $B^{2}_{n}\backslash T^{\prime}$ maximally until the NW, SW, SC, and SE strands are colored as Lemma 5.26 asks for. For the lower bound, we are performing a Type II replacement, where the quandle colorings at the endpoints of $T^{\prime}$ are as shown in Figure 32. ∎